OpenStudy (anonymous):

determine wheather 2,5,9,14 .... is arithmetic sequences.

OpenStudy (anonymous):

is 5-2 = 9-5?

OpenStudy (anonymous):

can i has a medal now? cause i think i hv the necessary koalafications (LOL) :D

OpenStudy (anonymous):

how do i write that out

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

no ur wrong it is arithmetic

OpenStudy (anonymous):

add 3 then add 4 then add 5

OpenStudy (anonymous):

2+3=5 5+4=9 9+5= 14

OpenStudy (anonymous):

oh wait yeah ur right @HabibiLove

OpenStudy (anonymous):

where did the 4 come from when you where doing 5+4 where did the 4 come from

OpenStudy (anonymous):

no wait im confused now.... i dont think it works that way....i was right in the beginning i think...

OpenStudy (anonymous):

that it isnt arthmetic sequence

OpenStudy (anonymous):

nvrmnd